# Improving Rectified Flow with Boundary Conditions

### v1

@@ 1 Introduction
Diffusion models have achieved significant success as a powerful class of generative models, demonstrating impressive results across various domains [37, 38, 12]. More recently, flow-based generative models have emerged as a compelling alternative for generative modeling [25, 20, 1, 10]. Within this family of flow-based models, Rectified Flow (RF) [25, 23], also known as Flow Matching [20] stands out by formulating generative modeling as solving an Ordinary Differential Equation (ODE) that directly transforms a noise distribution into the target data distribution.
@@
To address these limitations, we propose Boundary-enforced Rectified Flow Model (Boundary RF Model), a novel approach designed to explicitly enforce the crucial boundary velocities by incorporating constraints and informed parameterizations. Our method effectively rectifies the terminal behavior of the velocity field, ensuring adherence to boundary conditions, while preserving the representational power of state-of-the-art architectures like U-Net and Diffusion Transformers [29]. Moreover, Boundary RF Model provides a simple and readily implementable parameterization that enforces boundary conditions, easily integrable with existing Rectified Flow models.
@@ 2 Background
A standard way to solve RF ODE (1) is via the explicit Euler method. Let $10=t_{0}<t_{1}<\dots<t_{N}=$ 1 be $N$ timesteps in the interval $[0,1]$. The discretized trajectory $\{\tilde{{\bm{Z}}}_{t_{k}}\}$ evolves according to
@@ 3.2 Mask-based Boundary RF Model
$\min_{\theta}\int_{0}^{1}\mathbb{E}_{({\bm{X}}_{0},{\bm{X}}_{1})\sim\pi_{0}\times\pi_{1}}\left[\eta_{t}||{\bm{v}}({\bm{X}}_{t},t)-\dot{{\bm{X}}}_{t}||^{2}\right]\mathrm{d}t,$ (9)
@@
We will further explore alternative function choices and their empirical performance in Section 5.3.
@@ 3.4 Impact on Stochastic Sampling
As discussed, stochastic samplers for Rectified Flow inherently rely on the score function, which, via Tweedie’s formula (Eq. (5)), is inversely proportional to $(1-t)$. Consequently, if the boundary condition ${\bm{v}}(\mathbf{x},1)=x$ is not enforced, the score function estimate will become unbounded as $t\to 1$. This unboundedness destabilizes stochastic sampling, particularly when using larger step sizes (or $\sigma_{t}$) for faster generation. Such instability can manifest as over-smoothing and loss of detail in generated samples [14]. In contrast, Boundary RF Model, by enforcing ${\bm{v}}(\mathbf{x},1)=\mathbf{x}$, prevents the score function from diverging. This leads to a more stable score estimate, enabling robust stochastic sampling with larger $\sigma_{t}$.
@@ 4 Related Work
Diffusion models [12, 38, 37] have revolutionized the field of generative modeling by introducing iterative refinement for content generation. These models have demonstrated the capacity to generate samples from complex distributions across various data modalities, including images, videos, audio, and point clouds [6, 32, 13, 30, 3, 2]. A key characteristic of diffusion models is their reliance on iterative sampling over numerous steps, with the continuous limit being described by a Stochastic Differential Equation (SDE) process. In contrast, Rectified Flow  [25, 20, 1, 10], also known as Flow Matching, InterFlow, or IADB, presents an alternative approach to generative modeling based on Ordinary Differential Equations (ODEs). This ODE-based formulation is conceptually simpler and has proven to be successful in diverse practical applications, achieving performance comparable to or exceeding that of diffusion models in many cases [7, 8, 18, 31]. Recent research has increasingly recognized the conceptual connections and transferability between diffusion models and flow-based models [24, 9, 21]. Consequently, sampling approaches, training techniques, and other methodological details are being explored for interchangeable use between these frameworks [14]. In this work, we primarily focus on Rectified Flow due to its current efficacy and practical advantages in various applications. However, we note that the core concepts and contributions of our approach could also be potentially extended to diffusion models with minor adaptations.
@@
Existing approaches to improve the training of diffusion and flow models primarily focus on refining interpolation strategies and noise schedules [28, 19], or adjusting the weighting of the training objective across different timesteps [15, 7]. For example, Improved Denoising Diffusion Probabilistic Models (IDDPM) [28] proposed learning a more controlled noise schedule and variance. While these methods can be interpreted as implicitly improving model behavior near the data manifold, they do not explicitly modify the model architecture or directly address boundary conditions. In contrast, our work directly tackles a critical limitation of standard RF models by explicitly addressing and enforcing the theoretical boundary conditions. This complementary approach operates at the model parameterization level, offering a distinct perspective from existing training refinements.
@@ 5.2 Comparison with Vanilla RF Model
FID ()($\downarrow$)( ↓ ) IS ()($\uparrow$)( ↑ ) Precision ()($\uparrow$)( ↑ ) Recall ()($\uparrow$)( ↑ ) FID ()($\downarrow$)( ↓ ) IS ()($\uparrow$)( ↑ ) Precision ()($\uparrow$)( ↑ ) Recall ()($\uparrow$)( ↑ )
@@ 5.3 Ablation Studies
Right Boundary Left Boundary 2-Pass Model Euler-FID ()($\downarrow$)( ↓ ) Euler-IS ()($\uparrow$)( ↑ ) SDE-FID ()($\downarrow$)( ↓ ) SDE-IS ()($\uparrow$)( ↑ )
@@
The results, shown in Table 2, indicate that while both Mask-based Boundary RF Model outperform the vanilla RF model baseline in terms of FID, the one with double-boundary condition achieves slightly superior performance across most metrics than its counterpart with only right boundary, particularly in FID. The Subtraction-based Boundary RF Model provides a significant improvement over the baseline, demonstrating the effectiveness of enforcing at least the $t=1$ boundary condition. We note that Subtraction-based Boundary RF Model introduces approximately 2×2$\times$2 × increase in inference cost compared to the Mask-based Boundary RF Model, due to the extra forward pass for $t=1$.
@@ 5.5 Scalability to Large-Scale Models and High-Resolution Images
Method Euler-FID ()($\downarrow$)( ↓ ) Euler-IS ()($\uparrow$)( ↑ ) SDE-FID ()($\downarrow$)( ↓ ) SDE-IS ()($\uparrow$)( ↑ )
@@ 6 Conclusion
Future work could explore extending Boundary RF Model to larger models/higher resolutions or text-to-image/video generation task.
@@ A.1 Experiment Details
7:                $\hat{{\bm{Z}}}_{o}=\tilde{{\bm{Z}}}_{t_{i}}+({\bm{o}}-t_{i})\circ{\bm{v}}_{i},$   with    $o=\text{min}\big{(}t_{i+1}+c(t_{i+1}-t_{i}),~{}1\big{)}.$

### v2

@@ 1 Introduction
Diffusion models have achieved significant success as a powerful class of generative models, demonstrating impressive results across various domains [37, 38, 12]. More recently, flow-based generative models have emerged as a compelling alternative for generative modeling [25, 21, 1, 10]. Within this family of flow-based models, Rectified Flow (RF) [25, 23], also known as Flow Matching [21, 1] stands out by formulating generative modeling as solving an Ordinary Differential Equation (ODE) that transforms a noise distribution into the target data distribution.
@@
To address these limitations, we propose Boundary-enforced Rectified Flow Model (Boundary RF Model), a novel approach designed to explicitly enforce the boundary velocities by incorporating constraints and informed parameterizations. Our method effectively rectifies the terminal behavior of the velocity field, ensuring adherence to boundary conditions, while preserving the representational power of state-of-the-art architectures like U-Net and Diffusion Transformers [29]. Moreover, Boundary RF Model provides a simple and readily implementable parameterization that enforces boundary conditions, easily integrable with existing Rectified Flow models.
@@ 2 Background
A standard way to solve RF ODE (1) is via the explicit Euler method. Let $0=t_{0}<t_{1}<\dots<t_{N}=1$ be $N$ timesteps in the interval $[0,1]$. The discretized trajectory $\{\tilde{{\bm{Z}}}_{t_{k}}\}$ evolves according to
@@ 3.2 Mask-based Boundary RF Model
[…]
@@
We further explore alternative function choices and their empirical performance in Section 5.3.
@@ 3.4 Impact on Stochastic Sampling
As discussed, stochastic samplers for Rectified Flow inherently rely on the score function, which, via Tweedie’s formula (Eq. (5)), is inversely proportional to (1t)(1-t). Consequently, if the boundary condition ${\bm{v}}(\mathbf{x},1)=x$ is not enforced, the score function estimate will become unbounded as $t\to 1$. This unboundedness destabilizes stochastic sampling, particularly when using larger step sizes (or $\sigma_{t}$) for faster generation. Such instability can manifest as over-smoothing and loss of detail in generated samples [14]. In contrast, Boundary RF Model, by enforcing ${\bm{v}}(\mathbf{x},1)=\mathbf{x}$, prevents the score function from diverging. This leads to a more stable score estimate, enabling robust stochastic sampling with larger $\sigma_{t}$.
@@ 4 Related Work
Diffusion models [12, 38, 37] have revolutionized the field of generative modeling by introducing iterative refinement for content generation. These models have demonstrated the capacity to generate samples from complex distributions across various data modalities, including images, videos, audio, and point clouds [6, 32, 13, 30, 3, 2]. A key characteristic of diffusion models is their reliance on iterative sampling over numerous steps, with the continuous limit being described by a Stochastic Differential Equation (SDE) process. In contrast, Rectified Flow  [25, 21, 1, 10], also known as Flow Matching, InterFlow, or IADB, presents an alternative approach to generative modeling based on Ordinary Differential Equations (ODEs). This ODE-based formulation is conceptually simpler and has proven to be successful in diverse practical applications, achieving performance comparable to or exceeding that of diffusion models in many cases [7, 8, 19, 31]. Recent research has increasingly recognized the conceptual connections and transferability between diffusion models and flow-based models [24, 9, 22, 16]. Consequently, sampling approaches, training techniques, and other methodological details are being explored for interchangeable use between these frameworks [14, 24]. In this work, we primarily focus on Rectified Flow due to its current efficacy and practical advantages in various applications. However, we note that the core concepts and contributions of our approach could also be potentially extended to diffusion models with minor adaptations.
@@
Recent advances have improved flow and diffusion models’ training by adjusting the interpolation path or the noise schedule [28, 20, 16], or by explicitly re-weighting the loss across noise levels [15, 7, 16]. A series of analyses [24, 9, 22, 16] reveals that the noise schedule itself adds no expressiveness: different choices merely impose alternative weightings over noise levels. For instance, EDM [15] decouples the noise schedule from the auxiliary time variable, enabling sampling at arbitrary noise scales during inference. While other works [16, 24, 9] prove that any monotonic noise schedule reduces to a log-SNR-weighted loss. These methods, however, do not focus on the study of boundary conditions on RF. By contrast, our approach explicitly enforces the theoretical boundary conditions of the velocity network, remedying a shortcoming of standard RF parameterizations, offering a distinct perspective from existing training refinements.
@@ 5.2 Comparison with Vanilla RF Model
FID $(\downarrow)$ IS $(\uparrow)$ Precision $(\uparrow)$ Recall $(\uparrow)$ FID $(\downarrow)$ IS $(\uparrow)$ Precision $(\uparrow)$ Recall $(\uparrow)$
@@ 5.3 Ablation Studies
Right
Boundary
Left
Boundary
2-Pass
Model
Euler-FID $(\downarrow)$ Euler-IS $(\uparrow)$ SDE-FID $(\downarrow)$ SDE-IS $(\uparrow)$
@@
The results, shown in Table 2, indicate that while both Mask-based Boundary RF Model outperform the vanilla RF model baseline in terms of FID, the one with double-boundary condition achieves slightly superior performance across most metrics than its counterpart with only right boundary, particularly in FID. The Subtraction-based Boundary RF Model provides a significant improvement over the baseline, demonstrating the effectiveness of enforcing at least the $t=1$ boundary condition. We note that Subtraction-based Boundary RF Model introduces approximately $2\times$ increase in inference cost compared to the Mask-based Boundary RF Model, due to the extra forward pass for $t=1$.
@@ 5.5 Scalability to Large-Scale Models and High-Resolution Images
Method Euler-FID $(\downarrow)$ Euler-IS $(\uparrow)$ SDE-FID $(\downarrow)$ SDE-IS $(\uparrow)$
@@ 6 Conclusion
Future work could explore extending Boundary RF Model to larger models/higher resolutions on text-to-image/video generation task.
@@ A.1 Experiment Details
7:          $\hat{{\bm{Z}}}_{o}=\tilde{{\bm{Z}}}_{t_{i}}+({\bm{o}}-t_{i})\circ{\bm{v}}_{i},$   with    $\hskip 0.0pto=\text{min}\big(t_{i+1}+c(t_{i+1}-t_{i}),~1\big).$